# VGC-Bench: A Benchmark for Generalizing Across Diverse Team Strategies in Competitive Pokémon

### v2

@@ Abstract
Developing AI agents that can robustly adapt to dramatically different strategic landscapes without retraining is a central challenge for multi-agent learning. Pokémon Video Game Championships (VGC) is a domain with an extraordinarily large space of possible team configurations of approximately – far larger than those of Dota or StarCraft. The highly discrete, combinatorial nature of team building in Pokémon VGC causes optimal strategies to shift dramatically depending on both the team being piloted and the opponent’s team, making generalization uniquely challenging. To advance research on this problem, we introduce VGC-Bench: a benchmark that provides critical infrastructure, standardizes evaluation protocols, and supplies human-play datasets and a range of baselines – from large-language-model agents and behavior cloning to reinforcement learning and empirical game-theoretic methods such as self-play, fictitious play, and double oracle. In the restricted setting where an agent is trained and evaluated on a single-team configuration, our methods are able to win against a professional VGC competitor. We extensively evaluate all baseline methods over progressively larger team sets and find that even the best-performing algorithm in the single-team setting struggles at scaling up as team size grows. Thus, policy generalization across diverse team strategies remains an open challenge for the community. Our code is open sourced at https://github.com/cameronangliss/VGC-Bench.
1 Introduction
@@
Pokémon is the highest-grossing media franchise in the world, with an estimated total revenue exceeding $100 billion and a global player base numbering in the millions. In 2008, the Pokémon Company launched the Video Game Championships (VGC), a competitive series of tournaments featuring significant cash prizes and international prestige. The largest tournament to date, EUIC 2025, featured 1,257 competitors. Despite this popularity and competitive depth, to the best of our knowledge, no AI system has yet achieved human-level – let alone superhuman – performance in competitive Pokémon VGC battles.
@@
Competitive Pokémon presents two tightly-coupled challenges: team building and team usage. We estimate the size of the possible team configuration space in VGC to be on the order of , vastly exceeding the contextual complexity of other benchmark games such as Dota or StarCraft. Because of the discrete and combinatorial nature of team-building, optimal strategies can vary dramatically depending on the team compositions of both the player and opponent. This results in extreme context sensitivity, where even expert human players struggle to generalize strategies across matchups – even when their own team remains fixed, as is standard in VGC tournaments, but especially if trying to pilot many different team compositions. Some teams focus on controlling the weather; some focus on controlling the speed of the opponent; some focus on trapping the opponent’s Pokémon on the field. There is no well-defined maximum number of strategic paradigms with which a team can be built around, and some teams can be hybrid, playing to more than one of these powerful strategies at the same time.
This combination of a vast space of team configurations and highly diverse team strategies makes Pokémon an especially valuable testbed for generalization in AI systems. In the past, research has been conducted on Pokémon with heuristic and search-based agents [14, 12, 10], reinforcement learning [9, 20, 25, 15], team building [16], and predicting the winner at any state [4].
@@
In this work, we present VGC-Bench, a benchmark designed to evaluate AI generalization in Pokémon VGC. Our contributions include infrastructure for multi-agent learning and human-play data collection, a suite of competitive baselines, and robust evaluation tools for performance, generalization, exploitability, and human interaction. In the restricted setting where an agent is trained and evaluated on a single-team configuration, our methods are able to win against a professional VGC competitor. We then extend evaluation to broader multi-task settings involving diverse team configurations. While existing methods can learn competent policies in narrow conditions, they exhibit notable performance degradation as team diversity increases. These findings underscore the need for more generalizable learning algorithms in multi-agent, multi-task environments like VGC.
@@
Furthermore, as part of our work, we contributed significant open-source improvements to Poke-env [17], a widely used library that has supported prior research in Pokémon AI. Our contributions include full integration with the PettingZoo [21] multi-agent framework and extended support for VGC and doubles formats. We expect that these enhancements will enable and encourage more accessible and rigorous experimentation, further advancing AI research in this domain.
2 Problem Formulation
@@
VGC Mechanism: Pokémon VGC is a team-based competitive game with stochastic dynamics, simultaneous move selection, and a vast configuration space of possible teams. In the VGC format, each player assembles a team of 6 Pokémon, with each Pokémon configured by customized individual stats, up to 4 moves, a passive ability, a Tera type, and potentially a held item. VGC battles follow the doubles format, where players deploy up to 2 Pokémon on the field at a time. Recent tournaments adopt Open Team Sheets (OTS), a modification on the VGC format that reveals nearly all aspects about the opponent’s team – such as moves, items, abilities, and Tera types – while leaving the precise underlying stats concealed. Each match begins with a Team Preview phase, during which players simultaneously select 4 out of their 6 Pokémon to bring into the battle. The first two chosen are sent out at the start of the match, while the other two remain in reserve and can be switched in as the battle progresses. Once the battle begins, both players issue commands independently and simultaneously each turn, with Pokémon switching decision generally occurring before move choices. When all active Pokémon select moves, the execution order of actions is usually determined by their speed stats and speed ties are broken randomly. The objective is to knock out all of the opponent’s Pokémon by reducing their HP to zero before your own team is eliminated.
@@
Several factors make VGC particularly challenging for AI agents:
1.
@@
Combinatorial Team Configurations: The team configuration space in Pokémon VGC is extraordinarily large. With hundreds of Pokémon species, moves, items, and abilities, 19 possible Tera types per Pokémon, and a high-dimensional space of possible stat allocations, we estimate the total number of valid team configurations to be approximately . This combinatorial explosion far exceeds that of many other strategic games (see Table 1 for comparison).
2.
@@
3.
Partial Observability: Even with OTS, Pokémon VGC remains a partially observable game as the critical stats are not revealed. We estimate that the size of the information set – the set of all possible states given a partial observation of the game – is lower-bounded at approximately .
4.
@@
Formalization.
We model each Pokémon match as a two-player zero-sum partially observable stochastic game (POSG) with randomized team configurations. Let denote the finite set of legal team configurations, and let be the uniform distribution over configurations. At the beginning of each episode, players independently sample their teams , which jointly determine the instantiation of a POSG . Here, denotes the full battle state (including HP, status, weather, stat boosts, etc.), is the partial observation available to player , and is the set of available actions (select moves or switch Pokémon). is the stochastic transition function governed by Pokémon mechanics, and is the terminal reward function defined as if player 1 wins in terminal state and if player 2 wins. denotes the team configuration for both players.
Given policies and for the two players that map observation histories to action distributions, the expected return for player 1 across all team match-ups is defined as
where is the terminal state of trajectory . Since the game is zero-sum, the benchmark seeks a Nash equilibrium in expectation over configurations, which corresponds to the minimax-optimal pair of policies :
@@
3 Related Work
Progress in Game AI.
@@ 3 Related Work
Early AI milestones focused on perfect-information games like Chess and Go, where IBM’s Deep Blue and DeepMind’s AlphaGo achieved superhuman performance through massive search and pattern recognition [3, 19]. IBM’s Watson extended AI capabilities to natural-language trivia by winning Jeopardy! [5]. Poker introduced reasoning under partial observability and uncertainty, with Libratus defeating top humans in no-limit Texas Hold’em [2]. Recent work shifted to more complex multi-agent games: AlphaStar reached grandmaster in StarCraft II by population based reinforcement learning [24] and OpenAI Five mastered Dota 2 with limited hero drafting [1]. In contrast, Pokémon VGC offers a unique challenge that combines partial observability, stochasticity, simultaneous actions, and a combinatorially vast team configuration space larger than any prior benchmarked game (see Table 1).
| Game | Init. Config. Space | Partial Obs. | Stochastic | Simul. Actions |
| Chess | 2 | |||
| Go | 2 | |||
| Jeopardy! | 1 | |||
| Poker | ✓ | ✓ | ||
| StarCraft II | ✓ | ✓ | ✓ | |
| Dota 2 | ✓ | ✓ | ✓ | |
| Pokémon VGC | ✓ | ✓ | ✓ |
Existing Pokémon AI.
Early Pokémon AI research focused on handcrafted rules and search-based methods, such as those featured in the Showdown AI Competition[12], with agents using expectimax or minimax search[9], and heuristic bots designed to simulate human play [14]. Karten et al. [10] advanced this line with PokeChamp, combining minimax with LLMs for action sampling and opponent modeling. Deep reinforcement learning approaches have since demonstrated strong performance [8]: PPO agents trained via self-play[9], using Poke-env [20], or augmented with MCTS [25] have achieved expert-level play. However, most of the prior work focuses exclusively on the Single Battle format, where each player sends out one Pokémon at a time. This setting significantly simplifies the underlying decision space: in contrast, the official Pokémon VGC tournaments use Double Battle format, which involves two active Pokémon per side, leading to a combinatorial explosion in possible actions and interactions and more challenging learning. Moreover, prior evaluations are largely restricted to the Pokémon Showdown ladder, an online platform where opponents are often casual players of widely varying skill levels; by comparison, VGC-bench agents are directly evaluated against human experts and professional VGC players.
4 VGC-Bench
@@
This section introduces VGC-Bench, a comprehensive benchmark built on top of poke-env [17]. VGC-Bench provides users with (1) infrastructure for multi-agent learning and human-play data collection, (2) a diverse suite of baselines, and (3) robust evaluation tools for performance, generalization, exploitability, and human play.
4.1 Infrastructure
@@
We integrate poke-env with PettingZoo [21] to support parallelized population-based and PSRO[11]-style multi-agent learning for both players and support the VGC format. To support more controlled training settings, we introduce two toggles: one to exclude the team preview phase and another to disable mirror matches. (1) Disabling team preview skips that stage from the agent’s perspective, initializing each game with a random team preview decision – potentially encouraging policy exploration in any team preview situation. (2) Disabling mirror matches prevents an agent from facing the same team configuration during training, which is useful when evaluating a specific matchup between two distinct team configurations. However, we found that excluding the mirror match can result in reward hacking, where the agent learns to intentionally lose when playing one team and exploit the matchup when playing the other. Finally, we provide users with a pool of high-quality teams often used in real VGC competitions from VGCPastes. Teams are supplied as part of the environment configuration as plain text when a battle starts.
@@
We represent actions in the VGC format as a joint action space, since each turn in doubles matches requires an action for each of the two active Pokémon. The action space is enumerated in Table 9. Each Pokémon has 107 available actions, which captures the full space of switching in benched Pokémon and using moves, where moves involve which move is being picked, the intended target, and whether or not the Pokémon is terastallizing (or using another available gimmick). We also unify the team preview action of the battle with the action space for the rest of the battle by simply modeling team preview as two joint "switch-in" actions in a row, providing two Pokémon per joint action for a total of four Pokémon to make the team preview decision.
@@
To facilitate large-scale imitation learning from human players, we provide a parallel data collection pipeline for scraping and parsing Gen 9 VGC battle replays from Pokémon Showdown. With the current number of available replays from Pokémon Showdown, we are able to amass over 330,000 OTS-enabled VGC games. Due to the neutral perspective of battle logs and its purely transitional messages, reconstructing a trajectory of game states cannot be done perfectly. However, because we filter for only OTS-enabled battle logs, we can derive well-approximated states and actions from the logs by replaying the transitions through our environment.
4.2 Baseline Implementations

### v3

# VGC-Bench: Towards Mastering Diverse Team Strategies in Competitive Pokémon

@@ Abstract.
Developing AI agents that can robustly adapt to varying strategic landscapes without retraining is a central challenge in multi‑agent learning. Pokémon Video Game Championships (VGC) is a domain with a vast space of approximately team configurations, far larger than those of other games such as Chess, Go, Poker, StarCraft, or Dota. The combinatorial nature of team building in Pokémon VGC causes optimal strategies to vary substantially depending on both the controlled team and the opponent’s team, making generalization uniquely challenging. To advance research on this problem, we introduce VGC-Bench: a benchmark that provides critical infrastructure, standardizes evaluation protocols, and supplies a human-play dataset of over 700,000 battle logs and a range of baseline agents based on heuristics, large language models, behavior cloning, and multi-agent reinforcement learning with empirical game-theoretic methods such as self-play, fictitious play, and double oracle. In the restricted setting where an agent is trained and evaluated in a mirror match with a single team configuration, our methods can win against a professional VGC competitor. We repeat this training and evaluation with progressively larger team sets and find that as the number of teams increases, the best-performing algorithm in the single-team setting has worse performance and is more exploitable, but has improved generalization to unseen teams. Our code and dataset are open-sourced at https://github.com/cameronangliss/vgc-bench and https://huggingface.co/datasets/cameronangliss/vgc-battle-logs.
Key words and phrases:
Multi-Agent Learning, Reinforcement Learning, Benchmarking1. Introduction
@@
Pokémon is the highest-grossing media franchise in the world, with an estimated total revenue exceeding $100 billion and a global player base numbering in the millions. In 2008, the Pokémon Company launched the Video Game Championships (VGC), a competitive series of tournaments featuring significant cash prizes and international prestige. The largest tournament to date, EUIC 2025, featured 1,257 competitors. Despite this popularity and competitive depth, to the best of our knowledge, no AI system has yet achieved superhuman performance in competitive Pokémon VGC battles.
@@
Competitive Pokémon presents two tightly-coupled challenges: team building and team usage. Note that we only address team usage in this work; team building is left as an open challenge. We estimate the size of the team configuration space to be on the order of , vastly exceeding the configuration space of other benchmark games such as Dota or StarCraft. Because of the discrete and combinatorial nature of team-building, optimal strategies can vary dramatically depending on the team compositions of both the player and opponent. As a result, even expert human players struggle to generalize strategies across matchups – even when their own team remains fixed, as is standard in VGC tournaments, but especially if trying to pilot many different team compositions. Some teams focus on controlling the weather; some focus on controlling the speed of Pokémon on the field; some focus on preventing the opponent’s Pokémon from switching out. There is no well-defined maximum number of strategic paradigms with which a team can be built around, and many teams are hybrid, playing to more than one of these powerful strategies at the same time.
This combination of a vast space of team configurations and highly diverse team strategies makes Pokémon an especially valuable testbed for generalization in AI systems. In the past, research has been conducted on Pokémon with heuristic and search-based agents (panumate2016pokemon; lee2017showdown; karten2025pok), reinforcement learning (huang2019self; simoes2020competitive; wang2024winning; pleines2025pokemon), team building (simoes2023adversarial), and predicting the winner at any state (charde2019predicting). However, most of the prior work focuses exclusively on the Single Battle format, where each player sends out one Pokémon at a time. This setting significantly simplifies the underlying decision space: in contrast, the official Pokémon VGC tournaments use Double Battle format, which involves two active Pokémon per side, leading to a combinatorial explosion in possible actions and interactions and more challenging learning.
VGC-Bench captures the multi-agent multi-team dynamics with PettingZoo integration, provides human-play datasets and a range of baselines, and standardizes evaluation protocols.
@@
In this work, we present VGC-Bench, a benchmark designed to evaluate AI generalization in Pokémon VGC. Our contributions include infrastructure for multi-agent learning and human-play data collection, including a dataset of over 700,000 open team sheet VGC battle logs; a suite of competitive baselines, covering multi-agent RL, behavior cloning, using a language model, and heuristic agents; and robust evaluation tools for performance, generalization, exploitability, and human interaction. In the restricted setting where an agent is trained and evaluated on a single-team configuration, our methods are able to win against a professional VGC competitor. We then extend evaluation to broader multi-task settings involving diverse team configurations. While existing methods can learn competent policies in narrow conditions, they exhibit notable performance degradation as team diversity increases. These findings underscore the need for more generalizable learning algorithms in multi-agent, multi-task environments like VGC.
@@
Furthermore, as part of our work, we contributed significant open-source contributions to Poke-env (poke_env), a widely used library that has supported prior research in Pokémon AI. Our contributions include full integration with the PettingZoo (terry2021pettingzoo) multi-agent framework, extended support for VGC and doubles formats, and many bugfixes to correct battle tracking and more. We expect that these enhancements will enable and encourage more accessible and rigorous experimentation, further advancing AI research in this domain.
2. Problem Formulation
2.1. Game Mechanics
@@
Pokémon VGC is a team-based competitive game with stochastic dynamics, simultaneous move selection, and a vast configuration space of possible teams. In the VGC format, each player assembles a team of 6 Pokémon, with each Pokémon configured by customized individual stats, up to 4 moves, a passive ability, a Tera type, and potentially a held item. A Tera type is a gimmick mechanic specific to generation 9 of Pokémon which defines a type (i.e. grass, fire, water) that a Pokémon can overwrite their default types to for the rest of the battle. Players can only terastallize a Pokémon once per battle, and is commonly used defensively to avoid being super-effectively hit by an attack, but can also be used offensively to boost the damage of moves of the same type. The VGC format is entirely set in double battles, where players deploy up to 2 Pokémon on the field at a time. Recent tournaments adopt Open Team Sheets (OTS), a modification on the VGC format that reveals nearly all aspects about the opponent’s team – such as moves, items, abilities, and Tera types – while leaving the precise underlying stats concealed. Each match begins with a Team Preview phase, during which players simultaneously select 4 out of their 6 Pokémon to bring into the battle. The first two chosen are sent out at the start of the match, while the other two remain in reserve and can be switched in as the battle progresses. Once the battle begins, both players issue commands independently and simultaneously each turn, and only after the players have locked in their decisions do the battle mechanics determine the order of events for that turn. Generally, Pokémon switch out before moves are performed, and the execution order of moves is usually determined by the speed stats of the acting Pokémon, with speed ties broken randomly. The objective is to knock out all of the opponent’s Pokémon by reducing their HP to zero before your own team is eliminated.
@@
Several factors make VGC particularly challenging for AI agents (see Appendix A for calculation of values presented here):
@@
Combinatorial Team Configurations: The team configuration space in Pokémon VGC is large. With hundreds of Pokémon species, moves, items, and abilities, 19 possible Tera types per Pokémon, and a high-dimensional space of possible stat allocations, we estimate the total number of valid team configurations to be approximately . This combinatorial space far exceeds that of many other strategic games (see Table 1 for comparison).
@@
Partial Observability: Even with OTS, Pokémon VGC remains a partially observable game as the opponent’s stats are not revealed. We estimate that the size of the information set – the set of all possible states given a partial observation of the game – is lower-bounded at approximately .
@@
2.2. Formalization
We model each Pokémon battle as a two-player zero-sum partially observable stochastic game (POSG) with randomized team configurations. Let denote the finite set of legal team configurations and denote the uniform distribution over those team configurations. At the start of each episode, players independently and uniformly sample , which instantiates
Here, is the full battle state (HP, status, weather, stat boosts, etc.), is player ’s observation, the action set (switching, move selection, targeting, terastallization), the stochastic transition function, and the terminal reward for player 1.
Given policies mapping observations to action distributions, the distribution of trajectories initialized with team configurations can be defined as
The expected return for player 1 over all team draws is
Since the game is zero-sum, the equilibrium in expectation over configurations is
@@ 2.2. Formalization
Pokémon Showdown Gameplay. Closer Pokémon are on the agent’s side, and farther Pokémon are on the opponent’s side. 1) Pokémon’s health bar with percentage of current health remaining. 2) Current status of all party members, with solid colors for revealed, translucent colors for unrevealed, and greyed-out colors for fainted. 3) Effects on Pokémon, including boosts and status conditions. 4) Active side conditions and global fields/weather with a number of turns remaining. 5) Active Tera type being used by Pokémon.
3. Related Work
3.1. Progress in Game AI.
@@ 3. Related Work
Early AI milestones focused on perfect-information games like Chess and Go, where IBM’s Deep Blue and DeepMind’s AlphaGo achieved superhuman performance through massive search and pattern recognition (campbell2002deep; silver2016mastering). Poker introduced reasoning under partial observability and uncertainty, with Libratus defeating top humans in no-limit Texas Hold’em (brown2017libratus). Recent work shifted to more complex multi-agent games: AlphaStar reached grandmaster in StarCraft II by population-based reinforcement learning (vinyals2019grandmaster) and OpenAI Five mastered Dota 2 with limited hero drafting (berner2019dota). In contrast, Pokémon VGC offers a unique challenge that combines partial observability, stochasticity, simultaneous actions, and a combinatorially vast team configuration space larger than any prior benchmarked game (see Table 1).
| Game | Init. Config. Space |
| Chess | 2 |
| Go | 2 |
| Poker (10 players) | |
| StarCraft II | |
| Dota 2 | |
| Pokémon VGC |
3.2. Existing Pokémon AI
3.2.1. Heuristic and Search-Based Agents
Early Pokémon AI work relied on handcrafted rules and search. Lee and Togelius (2017) introduced the Showdown AI Competition, a benchmark based on a clone of Pokémon battles (lee2017showdown). They highlighted that Pokémon battles involve turn-based team combat with partial observability, a challenging setup uncommon in prior AI competitions. Huang and Lee (2019) note that before DRL, most agents used expectimax or minimax search with learned evaluations. Panumate and Iida (2016) built a simplified Generation-1 battle simulator and implemented four heuristic AIs (Random, Attack, Smart-Attack, Smart-Defense) to simulate and balance Pokémon gameplay (panumate2016pokemon). These simple bots modeled human behavior and were used to study “comfortable” game settings using game-refinement theory. In short, these studies demonstrate that search-based and rule-based methods can yield modest performance in Pokémon battles, but they often struggle with the game’s huge state space and complexity. Karten, Nguyen, and Jin (2025) made PokeChamp, an agent that plays Pokémon battles with a minimax search algorithm enhanced with an LLM (large language model). The LLM performs player action sampling, opponent modeling, and value function estimation, and achieves state-of-the-art LLM performance in the Gen 9 OU format (karten2025pok).
3.2.2. Reinforcement Learning Approaches
More recent work has applied deep RL to Pokémon. huang2019self trained a deep PPO agent via self-play for singles battles. Their approach requires no handcrafted simulator and succeeds on the nondeterministic, partially-observable Pokémon game, yielding a policy on par with state-of-the-art search agents and competitive with human ladder. simoes2020competitive similarly used deep RL on a Pokémon battling simulator (using the Poké-env gym interface) to learn competitive play. wang2024winning used DRL with PPO and augmented with parallel MCTS to train an AI agent to play Gen 4 random battles at an expert human level. pleines2025pokemon took this further by tackling Pokémon Red, an open-world RPG. They wrapped an emulator in an OpenAI Gym and trained a PPO agent to reach Cerulean City. Their experiments revealed that naive reward shaping can be exploited (e.g. agents skip intermediate challenges if badges give reward), and they underscore the exploration and multi-tasking challenges in Pokémon games. In summary, DRL (especially policy-gradient methods like PPO) has shown promise for learning Pokémon battle strategies from experience, achieving competitive performance without explicit simulators.
3.2.3. Team Building and Metagame Analysis
Building a strong team is crucial in Pokémon. simoes2023adversarial automated team construction under a metagame framework. They set up an adversarial model with two agents: a “team builder” that evolves six-Pokémon teams to maximize win-rate, and a “balancer” that adjusts base stats or move sets to encourage diversity in the meta. The team builders use evolutionary or linear optimization to exploit current metagames, while the balancer nudges them to try underused Pokémon. This adversarial setup iteratively yields balanced meta-teams.
3.2.4. Other Machine Learning Approaches
Some studies use supervised learning on Pokémon data. For example, charde2019predicting trained machine learning models to predict the winner of a Pokémon Showdown battle given the game state. This suggests that supervised classifiers can capture strategic patterns.
4. VGC-Bench
@@
This section introduces VGC-Bench, a comprehensive benchmark built on top of poke-env (poke_env). VGC-Bench provides users with (1) infrastructure for multi-agent learning and human-play data collection, (2) a diverse suite of baselines, and (3) robust evaluation tools for performance, generalization, exploitability, and human play.
4.1. Infrastructure
@@
We integrate poke-env with PettingZoo (terry2021pettingzoo) to support parallelized population-based and PSRO-style multi-agent learning (lanctot2017unified) for both players and support the VGC format. To support more controlled training settings, we introduce two toggles: one to exclude the team preview phase and another to disable mirror matches. Disabling team preview skips that stage from the agent’s perspective, initializing each game with a random team preview decision – potentially encouraging policy exploration in any team preview situation. Disabling mirror matches prevents an agent from facing the same team configuration during training, which is useful when evaluating a specific matchup between two distinct team configurations. Finally, we provide users with a team scraper that collects high-performing teams in real tournaments from VGCPastes (VGCPastes2025).
@@
We represent actions in the VGC format as a joint action space, since each turn in doubles matches requires an action for each of the two active Pokémon. The action space is enumerated in Table 6. Each Pokémon has 107 available actions, which captures the full space of switching in benched Pokémon and using moves, where moves involve which move is being picked, the intended target, and whether or not the Pokémon is terastallizing (or using another available gimmick). We also unify the team preview action of the battle with the action space for the rest of the battle by simply modeling team preview as two joint ”switch-in” actions in a row, providing two Pokémon per joint action for a total of four Pokémon to make the team preview decision.
@@
To facilitate large-scale imitation learning from human players, we provide a parallel data collection pipeline for scraping and parsing Gen 9 VGC battle replays from Pokémon Showdown. With the current number of available replays from Pokémon Showdown, we are able to amass over 700,000 OTS-enabled VGC games. Due to the neutral perspective of battle logs and its purely transitional messages, reconstructing a trajectory of game states cannot be done perfectly. However, because we filter for only OTS-enabled battle logs, we can derive well-approximated states and actions from the logs by replaying the transitions through our environment. Our log reader also has configurable filters to only read logs of players with ratings above a threshold and only the winner of the battle.
4.2. Baseline Implementations